OpenStudy (anonymous):

Which of the following is the equation of the circle with the center at the origin and r = 4?

OpenStudy (anonymous):

Equation of circle at center at origin is given by: \[\large \color{blue}{x^2 + y^2 = r^2}\] Here replace r =4.. Plug in the value and find the equation..

OpenStudy (anonymous):

\(r \) is radius here..

OpenStudy (anonymous):

\[r ^{2} = (x-h)^{2} + (y-k)^{2}\] where (h.k) = (0,0)